OpenStudy (anonymous):

If x, y, and z are positive numbers and 2x=3y=4z, then the value of x+y+z is how many times the value of x?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

First find the values of y and z in terms of x. Then insert them in x+y+z you will get \[\frac{ 13 }{ 6 }x\]This is to mean that x+y+z is 13/6 times grater than x.

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

x=1.5...

OpenStudy (anonymous):

y=4/3

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

Where you get this values?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

2x=3y

OpenStudy (anonymous):

x=1.5, right?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

Nope x=1.5y

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

6+4+3=13/6

OpenStudy (anonymous):

I plugged in my own values

OpenStudy (anonymous):

I found a common least multiple

OpenStudy (anonymous):

of x,y, and z

OpenStudy (anonymous):

Then we done right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

is that a final fantasy X boss you have there?